# Modeling Sampling Distributions of Test Statistics with Autograd

### v2

@@ 2.2 Flow-based Methods
by modeling the function $h:X\in\mathbb{R}^{d}\rightarrow Z\in\mathbb{R}^{d}$ with a sequence of bijections, where $g(*)$ is a tractable pdf and $|\det J|$ is the determinant of the Jacobian, $J$, of the transformation.
@@ 2.4 Conformal Inference
[…]
@@ 3.1 Modeling the cdf with ALFFI
Although the ALFFI algorithm yields valid confidence sets for the ON/OFF problem, it does not yield a sufficiently accurate approximation of the cdf, and hence of the pdf. This is shown in Fig. 1, where the empirical cdf [^2] Suppose that we have a sample of observations $\lambda_{1},...,\lambda_{n}$. The empirical cdf is defined by $F(\lambda)=\frac{1}{n}\sum_{i=1}^{n}\mathbb{I}\left(\lambda_{i}\leq\lambda\right)$, $F$, is displayed together with the approximated cdf $\hat{F}_{\text{ALFFI}}(\lambda\mid\mu,\nu)$. Taking the derivative of the approximated cdf yields the ALFFI approximation of the pdf, $\hat{f}_{\text{ALFFI}}(\lambda)=\frac{\partial\hat{F}_{\text{ALFFI}}}{\partial\lambda}$, displayed in Fig. 2. We observe that $\hat{f}_{\text{ALFFI}}(\lambda)$ displays sharp fluctuations and turning points, which was also confirmed by taking a numerical derivative of $\hat{F}_{\text{ALFFI}}$. This is because the actual slope of the cdf changes dramatically from point to nearby point, especially at low values of $\lambda$. It can also be seen that the histogrammed pmf of $\lambda$ at a particular $(\mu,\nu)$ point is a high frequency function, or one that is noisy and features very sharp peaks. The fluctuations seen in the pmf also depend on the choice of binning: finer binning shows larger fluctuations because this is a discrete distribution. This high-frequency nature of the pmf complicates our attempt to create a smooth modeling of the pmf with ALFFI.
@@ 5 Discussion and Conclusions
As can be seen in Eq. 5 conformal inference results in confidence sets that guarantee marginal coverage, $\mathbb{P}\left\{y_{n+1}\in\mathcal{C}_{\alpha}\left(x_{n+1}\right)\right\}\geq 1-\alpha$, that is, the coverage marginalized (integrated) over the population of which a dataset is a sample. That property is not at all controversial as coverage is a property of the population from which the dataset is presumed to have been drawn. But if each element $x$ of the dataset contains a subset that are parameters $\theta$ sampled from a prior $\pi_{\theta}$ then the marginalization is also with respect to this prior. When a prior is involved this form of coverage is weaker than the more sought after conditional coverage guarantee $\mathbb{P}\left\{y_{n+1}\in\mathcal{C}_{\alpha}\left(x_{n+1}\right)\mid\theta\right\}\geq 1-\alpha$. It was shown in [24] that such distribution-free conditional coverage is impossible to achieve with a finite sample, which has inspired studies that try to bridge the gap between marginal and conditional coverage [55]. In our case, $x=\{\theta,\lambda\}$ and we generate calibration sets at each $\theta$ point, thereby making our conformal confidence sets adaptive, that is, change with $\theta$. However, they are still not adaptive with respect to $\lambda$; the conformal confidence intervals are the same width at each $\lambda$ value. But it should be noted that this is not necessarily unusual. Consider, for example, the 68.3% confidence interval $[x-\sigma,x+\sigma]$ of a Gaussian pdf. The widths of these intervals are twice the standard deviation $\sigma$ and, therefore, independent of $x$.
@@
[…]
Acknowledgements This work is supported in part by the US Department of Energy under Award No. DE-SC0010102 and in part by the University Research Association (URA) Visiting Scholars Program Award No. 23-S-01. We are deeply thankful to Pushpalatha Bhat and Stephen Mrenna for their insightful feedback on this project and their patience. Special thanks to the Fermilab LHC Physics Center (LPC) for providing access to resources and facilities that were instrumental to the completion of this work.
@@ Appendix A ALFFI algorithm for modeling CDF for ON/OFF Example
1:estimated CDF ^$\widehat{\mathbb{C}}(\lambda\mid\mu,\nu)$ for all $\theta=\theta_{0}\in\Theta$
@@
16:Use $\mathcal{T^{\prime}}$ to learn the parameterized function ^$\widehat{\mathbb{C}}(\lambda\mid\mu,\nu):=\mathbb{E}[Z\mid\mu,\nu]$ via regression of $Z$ on $\{\mu,\nu,\lambda\}$ using mean square error as the loss function
17:return ^$\widehat{\mathbb{C}}(\lambda\mid\mu,\nu)$, which estimates $\mathbb{P}(\lambda_{i}\leq\lambda(N,M\mid\mu,\nu))$

### v3

@@ 2.2 Flow-based Methods
by modeling the function $h:X\in\mathbb{R}^{d}\rightarrow Z\in\mathbb{R}^{d}$ with a sequence of bijections, where g()g(*) is a tractable pdf and $|\det J|$ is the determinant of the Jacobian, $J$, of the transformation.
@@ 2.4 Conformal Inference
Conformal prediction (a.k.a. conformal inference) [21, 22, 23] is a general procedure for constructing such confidence sets/intervals for any predictive model (such as a neural network). These sets are valid (i.e., they satisfy Eq. 5) in finite samples without any assumptions about the distribution or the data other than the latter are exchangeable [21]. Split conformal prediction [24, 25] achieves this by splitting the $n$ points into a training set and a calibration set. A regression model $\hat{f}$ is fitted on the training set and then used to predict on the calibration set. Next a conformity score, for example, $s_{i}=|\hat{f}(x_{i})-y_{i}|$ is used to assess the agreement between the calibration’s response variable and the predicted value. Next, define $\hat{q}_{1-\alpha}$ to be the $\lceil(n+1)(1-\alpha)\rceil/n$ quantile of the scores $s_{1},...,s_{n}$, where $\lceil\cdot\rceil$ is the ceiling function. Finally, for a new test feature vector $x_{n+1}$ construct the conformal interval as
@@ 3.1 Modeling the cdf with ALFFI
Although the ALFFI algorithm yields valid confidence sets for the ON/OFF problem, it does not yield a sufficiently accurate approximation of the cdf, and hence of the pdf. This is shown in Fig. 1, where the empirical cdf [^2] Suppose that we have a sample of observations $\lambda_{1},...,\lambda_{n}$. The empirical cdf is defined by $(\lambda)=\frac{1}{n}\sum_{i=1}^{n}\mathbb{I}\left(\lambda_{i}\leq\lambda\right), FF$, is displayed together with the approximated cdf $\hat{F}_{\text{ALFFI}}(\lambda\mid\mu,\nu)$. Taking the derivative of the approximated cdf yields the ALFFI approximation of the pdf, $\hat{f}_{\text{ALFFI}}(\lambda)=\frac{\partial\hat{F}_{\text{ALFFI}}}{\partial\lambda}$, displayed in Fig. 2. We observe that $\hat{f}_{\text{ALFFI}}(\lambda)$ displays sharp fluctuations and turning points, which was also confirmed by taking a numerical derivative of $\hat{F}_{\text{ALFFI}}$. This is because the actual slope of the cdf changes dramatically from point to nearby point, especially at low values of $\lambda$. It can also be seen that the histogrammed pmf of $\lambda$ at a particular $(\mu,\nu)$ point is a high frequency function, or one that is noisy and features very sharp peaks. The fluctuations seen in the pmf also depend on the choice of binning: finer binning shows larger fluctuations because this is a discrete distribution. This high-frequency nature of the pmf complicates our attempt to create a smooth modeling of the pmf with ALFFI.
@@ 5 Discussion and Conclusions
As can be seen in Eq. 5 conformal inference results in confidence sets that guarantee marginal coverage, $\mathbb{P}\left\{y_{n+1}\in\mathcal{C}_{\alpha}\left(x_{n+1}\right)\right\}\geq 1-\alpha$, that is, the coverage marginalized (integrated) over the population of which a dataset is a sample. That property is not at all controversial as coverage is a property of the population from which the dataset is presumed to have been drawn. But if each element $x$ of the dataset contains a subset that are parameters $\theta$ sampled from a prior $\pi_{\theta}$ then the marginalization is also with respect to this prior. When a prior is involved this form of coverage is weaker than the more sought after conditional coverage guarantee $\mathbb{P}\left\{y_{n+1}\in\mathcal{C}_{\alpha}\left(x_{n+1}\right)\mid\theta\right\}\geq 1-\alpha$. It was shown in [24] that such distribution-free conditional coverage is impossible to achieve with a finite sample, which has inspired studies that try to bridge the gap between marginal and conditional coverage [55]. In our case, $x=\{\theta,\lambda\}$ and we generate calibration sets at each $\theta$ point, thereby making our conformal confidence sets adaptive, that is, change with $\theta$. However, they are still not adaptive with respect to $\lambda$; the conformal confidence intervals are the same width at each $\lambda$ value. But it should be noted that this is not necessarily unusual. Consider, for example, the 68.3% confidence interval $x-\sigma,x+\sigma]$ of a Gaussian pdf. The widths of these intervals are twice the standard deviation $\sigma$ and, therefore, independent of $x$.
@@ Appendix A ALFFI algorithm for modeling CDF for ON/OFF Example
1:estimated CDF $\widehat{\mathbb{C}}(\lambda\mid\mu,\nu)$ for all $\theta=\theta_{0}\in\Theta$
@@
16:Use $\mathcal{T^{\prime}}$ to learn the parameterized function $\widehat{\mathbb{C}}(\lambda\mid\mu,\nu):=\mathbb{E}[Z\mid\mu,\nu]$ via regression of $Z$ on $\{\mu,\nu,\lambda\}$ using mean square error as the loss function
17:return $\widehat{\mathbb{C}}(\lambda\mid\mu,\nu)$, which estimates $\mathbb{P}(\lambda_{i}\leq\lambda(N,M\mid\mu,\nu))$